OpenStudy (loser66):

How to find inverse of 11 mod 26? Please, help

OpenStudy (freckles):

\[ \text{ I think }11^{-1} \mod 26 \text{ is equvialent to solving } \\ 11x \equiv 1 \mod 26 \]

OpenStudy (ikram002p):

oh yeah @freckles is correct

OpenStudy (loser66):

oh, it is ikram, I am sorry friend, you changed your avatar. hehehe. I am trying to solve it now. @freckles Thanks for the tip :)

OpenStudy (loser66):

@ikram002p show me your way, please

OpenStudy (ikram002p):

ok so first way is trial but ur gotta be lucky , as seems 19 is the solution!

OpenStudy (freckles):

you can also do that euclidean thing

OpenStudy (loser66):

I have final on November 10. On my book, it says \(a^{-1}=a^{\phi (26)-1}\) I know how to find it, but it takes a long time to do. I am looking for the fastest way to find it out to save time on test.:)

OpenStudy (ikram002p):

this way is so fast with phi function

OpenStudy (loser66):

@freckles how to do with euclidean? @ikram002p no, it is not

OpenStudy (ikram002p):

but be careful its like this. a^phi(26)= 1 mod 26

OpenStudy (loser66):

phi (26) = 13, phi (26) -1 = 12 11^ 12 =? mod 26

OpenStudy (ikram002p):

well see phi(26)=12 (do u know how to get it ? )

OpenStudy (freckles):

\[11x-26k=1 \\ 26=11(2)+4 \\ 11=4(2)+3 \\ 4=3(1)+1 \\ \text{ now back through } \\ 4-3=1 \\ 4-[11-4(2)]=1 \\ 4-11+4(2)=1 \\ 4(3)-11=1 \\ 3(26-11(2)]-11=1 \\ 3(26)-11(6)-11=1 \\ 3(26)+11(-7)=1 \\ 11(-7)-26(-3)=1 \\ x= \equiv -7 \mod 26 =19\]

OpenStudy (loser66):

Yes, It is easier. Thanks @freckles

OpenStudy (ikram002p):

and @Loser66 here is the rule for any arbitrary number a,b such that gcd(a,b)=1 then \(\Large a^{\phi(b)}=1 \mod b\)

OpenStudy (ikram002p):

gcd(11,26)=1 then directly we can write \(\Large 11^{\phi(26)}=1 \mod 26 \\ \Large \text{ note that } \phi(26)=12 \\\Large 11^{12}=1 \mod 26 \)

OpenStudy (ikram002p):

ermm so for origin equation 11x=1 mod 26 x=11^11 mod 26 u can simplify it and u can use it like this only.

ganeshie8 (ganeshie8):

yeah that should work for \((a,n)=1\), from euler generalization of little fermat we have : \(1\equiv a^{\phi(n)}\pmod{n}\) multiply \(a^{-1}\) both sides

ganeshie8 (ganeshie8):

this is a pretty useless formula however reducing \(a^{\phi(n)-1}\) may not be any faster compared to other methods when the exponent or \(a\) is large enough

OpenStudy (ikram002p):

well its the fastest way compared to euclidean algorithm ;-), and yet it could be so fast as 11^6=-1 mod 26.

OpenStudy (loser66):

yes, that is the problem. And on test, I am not allowed to use calculator. :) one mistake can ruin my work.

ganeshie8 (ganeshie8):

trust your professor tha the gives human friendly numbers in the test

OpenStudy (loser66):

Euclidean method seems the best choice here. :) Thanks all for helping out.

OpenStudy (loser66):

I am rather well prepare for test than waiting for luck. :)

OpenStudy (ikram002p):

u can use a trick with phi(n) divisors idk its up to u @loser66. my teacher never came up with sensible numbers xD so cant hold on the trust only.

OpenStudy (loser66):

Thanks you guys.

OpenStudy (ikram002p):

=D

OpenStudy (ikram002p):

good luck though!

OpenStudy (loser66):

Thank you.